# Limit-case admissibility for positive infinite-dimensional systems

### v1

@@ 2.1 Transferring properties between ordered Banach spaces
As $D$ contains the constant one function $\operatorname{{\mathbbm{1}}}$, so $D$ is majorizing in $\widetilde{E}$. In particular, the canonical embeddings of $D$ into $E$ and $\widetilde{E}$ satisfy the assumptions of Theorem 2.10. Moreover, the bi-Laplace operator $A:=-\Delta^{2}$ with domain $E$ is the generator of an analytic uniformly eventually positive analytic $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $X$ by [13, Theorem 4.4], i.e., there exists $t_{0}\geq 0$ such that $T(t)\geq 0$ for all $t\geq t_{0}$. Furthermore, $\operatorname{dom}A^{n}\subseteq\widetilde{E}$ for sufficiently large $n\in\mathbb{N}$, so $T(t)X\subseteq D$ for all $t>0$. Theorem 2.10 thus ensures that there exists $c>0$ such that
$\left\lVert T(t)^{\prime}\right\rVert_{\operatorname{span}{E^{\prime}_{+}}\to L^{2}(\Omega)}\leq c\left\lVert T(t)\right\rVert_{L^{2}(\Omega)\to L^{\infty}(\Omega)}$
for all $t\geq t_{0}$.
Proof 2.11 (Proof of Theorem 2.10).
As $\tilde{E}_{+}$ is generating cone and $\tilde{J}(D)$ is majorizing in $\tilde{E}$, one can apply Proposition 2.1(c) to get $\tilde{J}^{\prime}(\tilde{E}^{\prime}_{+})=D^{\prime}_{+}$. On the other hand, the dual cone $\tilde{E}^{\prime}_{+}$ is also generating owing to the normality of the cone $\tilde{E}_{+}$ [4, Theorem 2.26] and thus $\tilde{J}^{\prime}\tilde{E}^{\prime}=D^{\prime}_{+}-D^{\prime}_{+}=\operatorname{span}D^{\prime}_{+}$. We now endow $\operatorname{span}D^{\prime}_{+}$ with the norm given in formula (1.2) (with $X_{+}$ replaced with $D^{\prime}_{+}$) and note that $\tilde{J}^{\prime}$ is continuous from $\tilde{E}^{\prime}$ to $\operatorname{span}D^{\prime}_{+}$ due to the closed graph theorem. Hence, the open mapping theorem implies that there exists a set $S\subseteq\tilde{E}^{\prime}$ that is norm-bounded by a number $\tilde{c}\geq 0$ such that $\tilde{J}^{\prime}S$ covers the unit ball in $\operatorname{span}D^{\prime}_{+}$. As a result,
@@
$\left\lVert(JT)^{\prime}\right\rVert_{\operatorname{span}\tilde{E}^{\prime}_{+}\to X^{\prime}}\leq\left\lVert T^{\prime}\right\rVert_{\operatorname{span}D^{\prime}_{+}\to X^{\prime}}\left\lVert J^{\prime}\right\rVert_{\operatorname{span}E^{\prime}_{+}\to\operatorname{span}D^{\prime}_{+}},$
which combined with the previous estimate completes the proof.
2.2 The order on $X_{-1}$
@@
$X_{-1,+}:=\overline{X_{+}}^{\left\lVert\mathord{\,\cdot\,}\right\rVert_{X_{-1}}},$
i.e., the norm closure of $X_{+}$ in $X_{-1}$, and consider the partial order that is induced by $X_{-1,+}$. This order on $X_{-1}$ was, for the case where $X$ is a Banach lattice, introduced in [10] in the context of perturbation theorems for positive semigroups. A number of fundamental properties of $X_{-1,+}$ – in particular that it is indeed a cone – are proved in [10, Remark 2.2 and Proposition 2.3]. We now show that the same properties remain true on ordered Banach spaces. Most arguments are similar, but we include the details for the convenience of the reader. Our proof of the equality $X_{-1,+}\cap(-X_{-1,+})=\{0\}$ is a bit different since the cone $X_{+}$ need not be normal in the following proposition.
@@
    One has $X_{+}=X_{-1,+}\cap X$. In other words, two vectors $x,y\in X$ satisfy $x\leq y$ in $X$ if and only if they satisfy the same inequality in $X_{-1}$.
@@
    For every $\lambda>s(A)$, the resolvent operator $R(\lambda,A_{-1})$ is positive from $X_{-1}$ to $X$.
Proof 2.13.
(c) Let $\lambda>s(A)=s(A_{-1})$. We prove that $R(\lambda,A_{-1})X_{-1,+}\subseteq X_{+}$ (this is precisely the claim of (c), but we are only allowed to speak of a positive operator from $X_{-1}$ to $X$ once we have proved (a)). So let $x\in X_{-1,+}$ and choose a sequence $(x_{n})\subseteq X_{+}$ that converges to $x$ in $X_{-1}$. Firstly, if $\lambda>\omega_{0}(A)$, then the operator $R(\lambda,A)$ is positive from $X$ to $X$ owing to the Laplace transform representation of the resolvent. Now, the positivity extends from the interval $(\omega_{0}(A),\infty)$ to the interval $(s(A),\infty)$ via the Taylor expansion of the resolvent and a connectedness argument; see for instance [23, Proposition 2.1(a)] for details. Therefore, $R(\lambda,A)x_{n}\in X_{+}$ for each index $n$. As $R(\lambda,A_{-1})$ is continuous from $X_{-1}$ to $X$, it follows that
@@
(b) Obviously, $X_{+}\subseteq X_{-1,+}\cap X$. To establish the converse inclusion, we let $x\in X_{-1,+}\cap X$. It follows from (c) that $\lambda R(\lambda,A)x=\lambda R(\lambda,A_{-1})x$ is in $X_{+}$ for all $\lambda>\max\{s(A),0\}$. Moreover, since $x$ is in $X$ one has $\lambda R(\lambda,A)x\to x$ in $X$ as $\lambda\to\infty$. As $X_{+}$ is closed in $X$, we even get that $x\in X_{+}$.
It is clear that the extrapolated semigroup $(T_{-1}(t))_{t\geq 0}$ on $X_{-1}$ leaves $X_{-1,+}$ invariant. Since $X_{-1}$ is, by the previous proposition, an ordered Banach space with respect to this cone, we can rephrase this by saying that the extrapolated semigroup is also positive.
Even if $X$ is a Banach lattice, the cone $X_{-1,+}$ is, in general, not generating in $X_{-1}$ [10, Examples 5.1 and 5.3]. In particular, $X_{-1}$ is usually not a Banach lattice. Nevertheless, for any $\lambda>s(A)$, we know from Proposition 2.12 that the operator $R(\lambda,A_{-1}):X_{-1}\to\operatorname{dom}A_{-1}=X$ is a continuous and positive linear bijection. This allows us to apply the results proved in Section 2.1 to $Z=X_{-1}$ in order to obtain various order properties of $X_{-1}$. First of all, Theorem 2.4 immediately implies the following (cf. Table 1):
@@
Let $(T(t))_{t\geq 0}$ be a positive $C_{0}$-semigroup on an ordered Banach space $X$. The following are true.
@@
    If the cone of $X$ is the face of the bidual cone, then the same is true for the cone of $X_{-1}$.
@@
As a consequence of Theorem 2.8, we get sufficient conditions for the cone $X_{+}$ to be a face of the cone $X_{-1,+}$.
@@
Proof 2.16.
Fix $\lambda>s(A)$. As pointed out before, $T=R(\lambda,A_{-1})$ is a continuous and positive linear bijection from $X_{-1}$ to $X$. Moreover, by definition $X$ is dense in $X_{-1}$ and the equality $X_{+}=X_{-1,+}\cap X$ from Proposition 2.12 implies the canonical embedding is bipositive. Therefore, defining $R_{n}:=nR(n,A_{-1})$ for $n\in\mathbb{N}$, Theorem 2.8 implies that $X_{+}$ is indeed a face of $X_{-1,+}$.
The following example shows that in general $X_{+}$ need not be a face in $X_{-1,+}$.
@@
On the space $E=\{f\in C[0,1]:f(0)=f(1)\}$, the operator
$\operatorname{dom}A:=\{f\in C^{1}[0,1]\in E:f^{\prime}(0)=f^{\prime}(1)\},\qquad f\mapsto f^{\prime}$
generates a positive periodic shift semigroup [41, Section A-I.2.5] for which the extrapolation space was shown in [10, Example 5.3] to be
@@
here $\mathcal{D}(0,1)$ is the space of test functions on $(0,1)$ and $\partial$ denotes the distributional derivative. The function
@@
lies in $X$ and satisfies $f-\partial f=\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}$. In particular, the characteristic function $\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}$ lies in $X_{-1,+}$ and satisfies $0\leq\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}\leq\operatorname{{\mathbbm{1}}}$. As $X_{+}$ contains $\operatorname{{\mathbbm{1}}}$ but not $\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}$, it follows that $X_{+}$ is not a face of $X_{-1,+}$.
We now present an application of Theorem 2.10 and follow it up with two examples involving elliptic operators. To state the result, let us first recall the definition of Favard spaces with index $\beta\in(-1,0]$. Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Banach space $X$, fix an arbitrary number $\omega$ that is strictly larger than the growth bound of the semigroup, and let $\beta\in(-1,0]$. For each $x\in X_{-1}$ one defines
@@
The space $F_{\beta}:=\{x\in X_{-1}:\,\left\lVert x\right\rVert_{\beta}<\infty\}$ is called the Favard space of order $\beta$ of the semigroup. It is a Banach space with respect to the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\beta}$; see, for instance, [17, Definition II.5.10 and II.5.11] for details.
@@
Let $X$ be a reflexive ordered Banach space, let $\varphi\in X^{\prime}$ be strictly positive, and let $(T(t))_{t\geq 0}$ be a positive and immediately differentiable $C_{0}$-semigroup. Assume that $T(t)^{\prime}X^{\prime}\subseteq(X^{\prime})_{\varphi}$ for each $t>0$ (where the definition of $(X^{\prime})_{\varphi}$ is given in (1.3)) and that $\operatorname{dom}A^{\prime}\cap(X^{\prime})_{\varphi}$ is majorizing in $(X^{\prime})_{\varphi}$.
@@
    $\left\lVert T_{-1}(t)\right\rVert_{\operatorname{span}X_{-1,+}\to X}\leq c\left\lVert T(t)^{\prime}\right\rVert_{X^{\prime}\to(X^{\prime})_{\varphi}}$
    for all $t>0$.
@@
    If $(T(t))_{t\geq 0}$ is additionally analytic and the dual semigroup satisfies the ultracontractivity type estimate $\left\lVert T(t)^{\prime}\right\rVert_{X^{\prime}\to(X^{\prime})_{\varphi}}\leq\widetilde{c}t^{-\alpha}$ for constants $\widetilde{c}\geq 0$ and $\alpha\in[0,1)$ and for all $t\in(0,1]$, then $\operatorname{span}X_{-1,+}$ is contained in the negative Favard space $F_{-\alpha}$ associated to $(T(t))_{t\geq 0}$.
Proof 2.19.
(a) We apply Theorem 2.10 to the following situation: choose the space $X$ from Theorem 2.10 as the space $X^{\prime}$ in the present theorem, let $E:=\operatorname{dom}A^{\prime}$, $\widetilde{E}:=(X^{\prime})_{\varphi}$, and $D:=E\cap\widetilde{E}$ which is an ordered Banach space with respect to the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{D}:=\left\lVert\mathord{\,\cdot\,}\right\rVert_{E}+\left\lVert\mathord{\,\cdot\,}\right\rVert_{\widetilde{E}}$. Let $J$ and $\widetilde{J}$ be the canonical embeddings of $D$ into $E$ and $\widetilde{E}$ and note that the range of $\widetilde{J}$ is majorizing in $\widetilde{E}$ by the assumptions of the present theorem. Fix $t>0$ and let $T:=T(t)^{\prime}$ – our assumptions imply that $\operatorname{Rg}T\subseteq D$. Hence, Theorem 2.10 yields a $c>0$, independent of $t$ such that
@@
Due to the reflexivity, one can replace the left-hand side of the estimate with the norm $\left\lVert T_{-1}(t)\right\rVert_{\operatorname{span}X_{-1,+}\to X}$, whence the claim.
(b) This is a consequence of (a) and a characterisation [17, Proposition II.5.13 and Definition II.5.11] of Favard spaces for analytic semigroups.
@@
holds. Consider the divergence form elliptic operator $A:L^{p}(\Omega)\supseteq\operatorname{dom}A\to L^{p}(\Omega)$, $u\mapsto\operatorname{div}(a\nabla u)$ with Neumann boundary conditions. This operator can be constructed using form methods in $L^{2}(\Omega)$ and then extrapolating to the $L^{p}$-scale, see [44, Section 4.1]. Let $q\in(1,\infty)$ be the Hölder conjugate of $p$. The operator $A$ generates a positive analytic $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $L^{p}(\Omega)$ [44, Corollary 4.3 and Theorem 1.52]. Further, if $\alpha:=\frac{d}{2q}<1$, then the following hold.
@@
    The span of the cone $L^{p}(\Omega)_{-1,+}$ is contained in the negative Favard space $F_{-\alpha}$.
  2. (b)
    The cone $L^{p}(\Omega)_{-1,+}$ consists precisely of all positive finite Borel measures on $\overline{\Omega}$.
Let us show how the claims (a) and (b) follow from the theory developed so far.
(a) The dual operator $A^{\prime}$ on the Banach space $L^{q}(\Omega)$ is also a divergence form elliptic operator with Neumann boundary conditions, but with the diffusion coefficient $a^{T}$. Therefore the semigroup $(T^{\prime}(t))_{t\geq 0}$ generated by $A^{\prime}$ maps $L^{q}$ into $L^{\infty}(\Omega)=\big{(}L^{q}(\Omega)\big{)}_{\operatorname{{\mathbbm{1}}}}$ and satisfies the ultracontractivity estimate
@@
(b) As $L^{p}(\Omega)$ is reflexive, the space $L^{p}(\Omega)_{-1}$ can be identified with $\operatorname{dom}(A^{\prime})^{\prime}$ as an ordered Banach space. So we only need to identify the positive linear functionals on $\operatorname{dom}A^{\prime}$. Since $\alpha<1$ it follows that $\operatorname{dom}A^{\prime}$ is contained in the space $C(\overline{\Omega})$ [43, Proposition 3.6]. Moreover, $\operatorname{dom}A^{\prime}$ is majorizing in $C(\overline{\Omega})$ since it contains the constant functions and it is dense in $C(\overline{\Omega})$ according to [43, Lemma 4.2]. Since the cone in $C(\overline{\Omega})$ has non-empty interior it follows that even the cone $\operatorname{dom}A^{\prime}_{+}$ is dense in the cone $C(\overline{\Omega})$. So by Proposition 2.1(b) the dual space $C(\overline{\Omega})^{\prime}$ embeds bipositively into the space $\operatorname{dom}(A^{\prime})^{\prime}\simeq L^{p}(\Omega)_{-1}$ and by part (c) of the same proposition, this embedding maps the cone of $C(\overline{\Omega})^{\prime}$ surjectively onto the cone of $\operatorname{dom}(A^{\prime})^{\prime}\simeq L^{p}(\Omega)_{-1}$. Since the positive cone in $C(\overline{\Omega})^{\prime}$ consists of all positive finite Borel measures on $\overline{\Omega}$, the claim follows.
@@
For elliptic operators with Dirichlet boundary conditions, the situation is a bit more subtle: in order to apply Theorem 2.18 one needs the semigroup to be intrinsically ultracontractive, which means that it not only maps $L^{1}$ into $L^{\infty}$ but also maps a weighted $L^{1}$-space – with the leading eigenfuntion $u$ of the operator as a weight – into the principal ideal generated by $u$. Such results are available in the literature in the case that the domain and the coefficients of the differential operator are sufficiently smooth, see for instance, [14, Theorem 4.6.2]. To get an estimate from $L^{p}$ into the principal ideal generated by $u$ we need $p$ to be at least $2$ – this is why the result in the following example can only be applied if the dimension of the spatial domain is low enough.
@@
Example 2.21 ($X_{-1}$ for elliptic operators with Dirichlet boundary conditions).

### v4

@@ 2.1. Transferring properties between (pre-)ordered Banach spaces
The bi-Laplace operator $A:=-\Delta^{2}$ with domain $E$ generates a uniformly eventually positive analytic $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $X$ by [16, Theorem 4.4], i.e., there exists $t_{0}\geq 0$ such that $T(t)\geq 0$ for all $t\geq t_{0}$. Furthermore, $\operatorname{dom}A^{n}\subseteq\widetilde{E}$ for large $n\in\mathbb{N}$, so $T(t)X\subseteq D$ for all $t>0$. As in (a), it can be verified that the assumptions of Theorem 2.6 are fulfilled and whence there exists $c>0$ such that for all $t\geq t_{0}$,
$\left\lVert T(t)^{\prime}\right\rVert_{\operatorname{span}{E^{\prime}_{+}}\to L^{2}}\leq c\left\lVert T(t)\right\rVert_{L^{2}\to L^{\infty}}.$
Proof of Theorem 2.6.
As $\tilde{E}_{+}$ is generating in $\tilde{E}$ and $\tilde{J}(D)$ is majorizing in $\tilde{E}$, one can apply Proposition 2.1(a) and (c) to get $\tilde{J}^{\prime}(\tilde{E}^{\prime}_{+})=D^{\prime}_{+}$. On the other hand, the dual wedge $\tilde{E}^{\prime}_{+}$ is also generating owing to the normality of the cone $\tilde{E}_{+}$ [4, Theorem 2.26] and thus $\tilde{J}^{\prime}\tilde{E}^{\prime}=D^{\prime}_{+}-D^{\prime}_{+}=\operatorname{span}D^{\prime}_{+}$. Note that $\tilde{J}^{\prime}$ is continuous from $\tilde{E}^{\prime}$ to $\operatorname{span}D^{\prime}_{+}$ due to the closed graph theorem; here $\operatorname{span}D^{\prime}_{+}$ is endowed with the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{D^{\prime}_{+}-D^{\prime}_{+}}$. Hence, the open mapping theorem implies that there exists a set $S\subseteq\tilde{E}^{\prime}$ that is norm-bounded by some $\tilde{c}\geq 0$ such that $\tilde{J}^{\prime}S$ covers the unit ball in $\operatorname{span}D^{\prime}_{+}$. So,
@@
$\left\lVert(JT)^{\prime}\right\rVert_{\operatorname{span}E^{\prime}_{+}\to X^{\prime}}\leq\left\lVert T^{\prime}\right\rVert_{\operatorname{span}D^{\prime}_{+}\to X^{\prime}}\left\lVert J^{\prime}\right\rVert_{\operatorname{span}E^{\prime}_{+}\to\operatorname{span}D^{\prime}_{+}},$
which completes the proof when combined with the previous estimate. ∎
2.2. The order on $X_{-1}$
@@
$X_{-1,+}:=\overline{X_{+}}^{\left\lVert\mathord{\,\cdot\,}\right\rVert_{-1}},$
i.e., the norm closure of $X_{+}$ in $X_{-1}$, and consider the partial order that is induced by $X_{-1,+}$. This order on $X_{-1}$ was, for the case where $X$ is a Banach lattice, introduced in [13] in the context of perturbation theorems for positive semigroups. A number of fundamental properties of $X_{-1,+}$ – in particular that it is indeed a cone – are proved in [13, Remark 2.2 and Proposition 2.3]. We now show that the same properties remain true on ordered Banach spaces. Most arguments are similar, but we include the details for the convenience of the reader. Our proof of the equality $X_{-1,+}\cap(-X_{-1,+})=\{0\}$ is a bit different since the cone $X_{+}$ need not be normal in the following proposition.
@@
    The canonical embedding $X\hookrightarrow X_{-1}$ is bipositive, i.e., $X_{+}=X_{-1,+}\cap X$.
@@
    For any $\lambda>s(A)$, the resolvent $R(\lambda,A_{-1})$ is positive from $X_{-1}$ to $X$.
Proof.
(c) Let $\lambda>s(A)=s(A_{-1})$. We need to prove that $R(\lambda,A_{-1})X_{-1,+}\subseteq X_{+}$. So let $x\in X_{-1,+}$ and choose a sequence $(x_{n})\subseteq X_{+}$ that converges to $x$ in $X_{-1}$. Firstly, if $\lambda>\omega_{0}(A)$, then the operator $R(\lambda,A)$ is positive from $X$ to $X$ owing to the Laplace transform representation of the resolvent. Now, the positivity extends from the interval $(\omega_{0}(A),\infty)$ to the interval $(s(A),\infty)$ via the Taylor expansion of the resolvent and a connectedness argument; see for instance [26, Proposition 2.1(a)] for details. Therefore, $R(\lambda,A)x_{n}\in X_{+}$ for each index $n$. As $R(\lambda,A_{-1})$ is continuous from $X_{-1}$ to $X$, it follows that
@@
(b) Obviously, $X_{+}\subseteq X_{-1,+}\cap X$. To establish the converse inclusion, we let $x\in X_{-1,+}\cap X$. It follows from (c) that $\lambda R(\lambda,A)x=\lambda R(\lambda,A_{-1})x\in X_{+}$ for all $\lambda>\max\{s(A),0\}$. Moreover, since $x$ is in $X$ one has $\lambda R(\lambda,A)x\to x$ in $X$ as $\lambda\to\infty$. As $X_{+}$ is closed in $X$, we even get that $x\in X_{+}$. ∎
Clearly, the extrapolated semigroup $(T_{-1}(t))_{t\geq 0}$ on $X_{-1}$ leaves $X_{-1,+}$ invariant. Since $X_{-1}$ is, by the previous proposition, an ordered Banach space with respect to this cone, we can rephrase this by saying that the extrapolated semigroup is positive.
Even if $X$ is a Banach lattice, the cone $X_{-1,+}$ need not be generating in $X_{-1}$ [13, Examples 5.1 and 5.3]. In particular, $X_{-1}$ is usually not a Banach lattice. One can show that the span of $X_{-1,+}$ is often a Banach lattice, though [10, Section 4], but we will not use the observation in what follows. Nevertheless, for any $\lambda>s(A)$, we know from Proposition 2.8 that the bijection $R(\lambda,A_{-1})\in\mathcal{L}(X_{-1},X)_{+}$. Hence we can apply the results proved in Section 2.1 to $Z=X_{-1}$ to obtain various order properties of $X_{-1}$. Firstly, Theorem 2.3 readily implies the following result (cf. Table 1). Thereafter, we obtain sufficient conditions for the cone $X_{+}$ to be a face of the cone $X_{-1,+}$.
@@
Let $(T(t))_{t\geq 0}$ be a positive $C_{0}$-semigroup on an ordered Banach space $X$.
@@
    If the cone of $X$ is a face of the bidual wedge, then the same is true for the cone of $X_{-1}$.
@@
Proof.
By definition, $X$ is dense in $X_{-1}$ and according to Proposition 2.8(b) the canonical embedding is bipositive. Therefore, defining $T=R(\lambda,A_{-1})$ for fixed $\lambda>s(A)$ and $R_{n}:=nR(n,A_{-1})$ for sufficiently large $n\in\mathbb{N}$, Theorem 2.5 implies that $X_{+}$ is indeed a face of $X_{-1,+}$. ∎
The following example shows that Corollary 2.10 fails if $X_{+}$ is not a face of $X^{\prime\prime}_{+}$.
@@
On the space $X=\{f\in C[0,1]:f(0)=f(1)\}$, the operator
$\operatorname{dom}A:=\{f\in C^{1}[0,1]\in X:f^{\prime}(0)=f^{\prime}(1)\},\qquad f\mapsto f^{\prime}$
generates a positive periodic shift semigroup [47, Section A-I.2.5] for which the extrapolation space was shown in [13, Example 5.3] to be
@@
here $\mathcal{D}(0,1)$ is the space of test functions on $(0,1)$ and $\partial$ is the distributional derivative. The function
@@
lies in $X$ and satisfies $f-\partial f=\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}$. In particular, $\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}\in X_{-1}$ and $0\leq\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}\leq\operatorname{{\mathbbm{1}}}$. As $X_{+}$ contains $\operatorname{{\mathbbm{1}}}$ but not $\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}$, we conclude that $X_{+}$ is not a face of $X_{-1,+}$.
We now present an application of Theorem 2.6 and follow it up with two examples involving elliptic operators. To state the result, let us first recall the definition of Favard spaces with index $\beta\in(-1,0]$. Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Banach space $X$ with generator $A$ and fix $\omega>\omega_{0}(A)$. For each $x\in X_{-1}$, one defines
@@
The space $F_{\beta}:=\{x\in X_{-1}:\,\left\lVert x\right\rVert_{\beta}<\infty\}$ is called the Favard space of order $\beta$ of the semigroup. It is a Banach space with respect to the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\beta}$; see, for instance, [20, Definitions II.5.10 and II.5.11] for details.
@@
Let $(T(t))_{t\geq 0}$ be a positive and immediately differentiable $C_{0}$-semigroup on a reflexive ordered Banach space $X$ whose cone $X_{+}$ is generating. Let $\varphi\in X^{\prime}$ be strictly positive. Assume that $T(t)^{\prime}X^{\prime}\subseteq(X^{\prime})_{\varphi}$ for all $t>0$ and $\operatorname{dom}A^{\prime}\cap(X^{\prime})_{\varphi}$ is majorizing in $(X^{\prime})_{\varphi}$.
@@
    $\left\lVert T_{-1}(t)\right\rVert_{\operatorname{span}X_{-1,+}\to X}\leq c\left\lVert T(t)^{\prime}\right\rVert_{X^{\prime}\to(X^{\prime})_{\varphi}}\qquad(t>0).$
@@
    If $(T(t))_{t\geq 0}$ is analytic and the dual semigroup satisfies an ultracontractivity type estimate $\left\lVert T(t)^{\prime}\right\rVert_{X^{\prime}\to(X^{\prime})_{\varphi}}\leq\widetilde{c}t^{-\alpha}$ for some $\widetilde{c}\geq 0$, $\alpha\in[0,1)$, and all $t\in(0,1]$, then $\operatorname{span}X_{-1,+}$ is contained in the Favard space $F_{-\alpha}$.
For the definition of $(X^{\prime})_{\varphi}$, we refer to (1.3). We point out that under the assumptions of Theorem 2.12, the extrapolation semigroup $(T_{-1}(t))_{t\geq 0}$ is also immediately differentiable and hence maps $X_{-1}$ to $\operatorname{dom}A_{-1}=X$ for all $t>0$.
Proof of Theorem 2.12.
(a) We apply Theorem 2.6 to the following situation: choose the space $X$ from Theorem 2.6 as the space $X^{\prime}$ in the present theorem, let $E:=\operatorname{dom}A^{\prime}$ and $\widetilde{E}:=(X^{\prime})_{\varphi}$, Note that as the cone in $X$ is generating, the cone in $X^{\prime}$ is normal [4, Theorem 2.42] and hence $\widetilde{E}=(X^{\prime})_{\varphi}$ is an ordered Banach space with a normal cone with respect to the gauge norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\varphi}$ [4, Theorems 2.60 and 2.63]. Moreover, the cone in $\widetilde{E}$ can easily be checked to be generating. Finally, choose $D:=E\cap\widetilde{E}$ which is an ordered Banach space with respect to the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{D}:=\left\lVert\mathord{\,\cdot\,}\right\rVert_{E}+\left\lVert\mathord{\,\cdot\,}\right\rVert_{\widetilde{E}}$. Let $J$ and $\widetilde{J}$ be the canonical embeddings of $D$ into $E$ and $\widetilde{E}$ and note that the range of $\widetilde{J}$ is majorizing in $\widetilde{E}$ by the assumptions of the present theorem. Fix $t>0$ and let $T:=T(t)^{\prime}$ – our assumptions imply that $\operatorname{Rg}T\subseteq D$. Hence, Theorem 2.6 yields a $c>0$, independent of $t$, such that
@@
Due to the reflexivity, $X_{-1}$ can be identified with $\operatorname{dom}(A^{\prime})^{\prime}$ as an ordered Banach space. Therefore, one can replace the left-hand side of the previous estimate with $\left\lVert T_{-1}(t)\right\rVert_{\operatorname{span}X_{-1,+}\to X}$ and doing so yields the claim.
(b) This is a consequence of (a) and a characterisation [20, Proposition II.5.13 and Definition II.5.11] of Favard spaces for analytic semigroups. ∎
@@
holds. Consider the divergence form elliptic operator $A:L^{p}(\Omega)\supseteq\operatorname{dom}A\to L^{p}(\Omega)$, $u\mapsto\operatorname{div}(a\nabla u)$ with Neumann boundary conditions. This operator can be constructed using form methods in $L^{2}(\Omega)$ and then extrapolating to the $L^{p}$-scale, see [50, Section 4.1]. The operator $A$ generates a positive analytic $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $L^{p}(\Omega)$ [50, Corollary 4.3 and Theorem 1.52]. Let $q\in(1,\infty)$ be the Hölder conjugate of $p$ and suppose $\alpha:=\frac{d}{2q}<1$. Then the following hold.
@@
    The span of the cone $L^{p}(\Omega)_{-1,+}$ is contained in the Favard space $F_{-\alpha}$. Indeed, the dual operator $A^{\prime}$ on the Banach space $L^{q}(\Omega)$ is also a divergence form elliptic operator with Neumann boundary conditions, but with the diffusion coefficient $a^{T}$. Therefore the semigroup $(T^{\prime}(t))_{t\geq 0}$ generated by $A^{\prime}$ maps $L^{q}$ into $L^{\infty}(\Omega)=\big{(}L^{q}(\Omega)\big{)}_{\operatorname{{\mathbbm{1}}}}$ and satisfies the ultracontractivity estimate
@@
    The cone $L^{p}(\Omega)_{-1,+}$ is the set of all positive finite Borel measures on $\overline{\Omega}$. Indeed, as $L^{p}(\Omega)$ is reflexive, the space $L^{p}(\Omega)_{-1}$ can be identified with $\operatorname{dom}(A^{\prime})^{\prime}$ as an ordered Banach space. So we only need to identify the positive linear functionals on $\operatorname{dom}A^{\prime}$. Since $\alpha<1$ it follows that $\operatorname{dom}A^{\prime}$ is contained in the space $C(\overline{\Omega})$ [49, Proposition 3.6]. Moreover, $\operatorname{dom}A^{\prime}$ is majorizing in $C(\overline{\Omega})$ since it contains the constant functions and is dense in $C(\overline{\Omega})$ by [49, Lemma 4.2]. Since the cone in $C(\overline{\Omega})$ has non-empty interior it follows that even the cone $(\operatorname{dom}A^{\prime})_{+}$ is dense in the cone $C(\overline{\Omega})_{+}$, so by Proposition 2.1(b) the dual space $C(\overline{\Omega})^{\prime}$ embeds bipositively into the space $\operatorname{dom}(A^{\prime})^{\prime}\simeq L^{p}(\Omega)_{-1}$ and by part (c) of the same proposition, this embedding maps $C(\overline{\Omega})^{\prime}_{+}$ surjectively onto the cone of $\operatorname{dom}(A^{\prime})^{\prime}\simeq L^{p}(\Omega)_{-1}$. As the positive cone in $C(\overline{\Omega})^{\prime}$ consists of all positive finite Borel measures on $\overline{\Omega}$, the claim follows.
@@
For elliptic operators with Dirichlet boundary conditions, the situation is more subtle: in order to apply Theorem 2.12 one needs the semigroup to be intrinsically ultracontractive, meaning that it not only maps $L^{1}$ into $L^{\infty}$ but also maps a weighted $L^{1}$-space – with the leading eigenfuntion $u$ of the operator as a weight – into the principal ideal generated by $u$. Such results exist in the literature in the case that the domain and the coefficients of the differential operator are sufficiently smooth, see for instance, [17, Theorem 4.6.2]. However, to get an estimate from $L^{p}$ into the principal ideal generated by $u$ we also need $p$ to be at least $2$ – this is why the result in the following example can only be applied if the spatial domain is an interval.
@@
Example 2.14 ($X_{-1}$ for elliptic operators with Dirichlet boundary conditions on intervals).